• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 08:34:09 on March 05, 2012

    Entry number 363774

    keyword=Compton electron detector runs

    Compton tune was good with photon. Though I only got like 30 minutes for
    edet right before access.
    I had to play with threholds but could not see the Compton edge.
    I still need to look through the data I took to try to optimize
    thresholds. But so far I have not seen good compton events from electron
    detector ( no change laser on /off ).
    24961 HV ramping
    24962 good run 25 k events
    24963 Good run threshold too
    24964 changing threshold from 64 to 7*16
    24965 7*16 thres
    24966 edet moving
    24967 edet moving to 7 mm
    24968 edet 7 mm threshold 0x70 1 plane trigger
    24969 2 plane trigger 0x40
    24670 same
    24671 Left thresh A C D to 0x4f
    24675 Left thresh D 0x4f
    24676 same
    24677 same
    24978 same 64 channels

    A copy of this log entry has been emailed to: nanda@jlab.org